a 6 kg bomb explode into 3 equal pieces  P Q and R. if P flies with a speed of 30m/s and  Q with 40 m/s making angle 90 degree with the direction of P. find the angle between P and R?
 

Grade:12th pass

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
Let us take P moving along y axis, Q along x axis and R making an angle alpha with -x axis and is in the third quadrant.
Hence, by conservation of momentum, we get:
Along y axis:
30\times m=mvsin\alpha\Rightarrow vsin\alpha=30\quad...(i)
Along x axis:
40\times m=mvcos\alpha\Rightarrow vcos\alpha=40\quad...(ii)
Hence,
tan\alpha=\frac{3}{4}\Rightarrow \alpha=tan^{-1}\frac{3}{4}=37^{\circ}
So, the required angle is equal to90^{\circ}+37^{\circ}=127^{\circ}between P and R.
Regards